OpenStudy (anonymous):

solve by using elimination.express answer as an ordered pair 2x-5y=-7 5x-3y=11

OpenStudy (michele_laino):

hint: we can multiply the first equation by 5, so the new first equation is: 10x-25y=-35 next we can multiply the second equation by 2, so we get: 10x-6y=22 then we have this equivalent system: \[\Large \left\{ \begin{gathered} 10x - 25y = - 35 \hfill \\ 10x - 6y = 22 \hfill \\ \end{gathered} \right.\] now, please subtract the second equation, from the first one, what equation do you get?

OpenStudy (anonymous):

10x-19y=-13

OpenStudy (michele_laino):

hint: if we do that subtraction, we can write this: \[\Large \left( {10x - 25y} \right) - \left( {10x - 6y} \right) = - 35 - 22\] please simplify

OpenStudy (anonymous):

im confused im i supposed to subtract -35 and 22

OpenStudy (michele_laino):

yes!: -35-22=-57

OpenStudy (michele_laino):

now, please simplify the left side

OpenStudy (anonymous):

25y and 6y = -31

OpenStudy (michele_laino):

hint: \[\Large \begin{gathered} \left( {10x - 25y} \right) - \left( {10x - 6y} \right) = \hfill \\ \hfill \\ = 10x - 25y - 10x + 6y = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

-15-4=-19

OpenStudy (michele_laino):

what are -15 and -4 ?

OpenStudy (anonymous):

you said 10x-25y

OpenStudy (michele_laino):

here is the right step: 

OpenStudy (anonymous):

oooo

OpenStudy (michele_laino):

so we get this equation: \[\Large - 19y = - 57\] please solve for y

OpenStudy (anonymous):

3

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

now, please replace y with 3 into the first equation of the original system, what do you get?

OpenStudy (anonymous):

15

OpenStudy (michele_laino):

yes! and Iwe gaet the subsequent equation: \[\Large 2x - 15 = - 7\] please solve for x

OpenStudy (michele_laino):

we*

OpenStudy (anonymous):

4

OpenStudy (michele_laino):

correct! so the solution of your system is: x=4 y=3

OpenStudy (anonymous):

thank you for helping me i understand it now a little

OpenStudy (michele_laino):

:)